OpenStudy (anonymous):

Internal telephone numbers in a university are composed of 5 digits. The first two digits can form any integer between (and including) 61 and 75, the third digit is an integer between 1 and 9 and the last two each take any integer value. Assume that it is disallowed to have a phone number with last three digits being 100, 101, or 102. In such case, the total number of possible internal telephone numbers that start with the digit 7 is.....

OpenStudy (anonymous):

is it 4500 - 15 = 4485

OpenStudy (amistre64):

75 61 --- 14+1 = 15 ways 1 to 9 = 9 ways (if we can count 1 and 9) 1000 ways - 3 = 997 ways for the ending 15*9*997 is what it looks like to me

OpenStudy (amistre64):

is 134,595 an option?

OpenStudy (anonymous):

the first digit starts with a 7 mate, so its 5 possible choices times by 9 possible choices times by 100 possible choices for the last 2 digits including "00", since i have to avoid 100,101,102, i got to take of 15 possibilities so its 4500-15 = 4485

OpenStudy (amistre64):

yeah, start with a 7 ... to many numbers to sort thru lol 70 71 72 73 74 75 = 6 ways then 6.9.997 is what I would see then

OpenStudy (amistre64):

the last 3 digits can be from 000 to 999 (1000), excluding the 3 ways mentioned = 997 ways

OpenStudy (amistre64):

61 to 75 = 15 ways, exclude the 6s and we get 6 ways to start

OpenStudy (amistre64):

53,838 an option?

OpenStudy (anonymous):

so its 5400 - 18 = 5382, its not multiple choice =\

OpenStudy (amistre64):

wait, im still reading the question a little off. try this:

OpenStudy (amistre64):

composed of 5 digits. first two digits, starts with a 7: 7 __ 6 ways 3rd digit between 1 and 9: 9 ways last two digits each take any integer value. Assume that it is disallowed to have a phone number with last three digits being 100, 101, or 102

OpenStudy (amistre64):

100 to 999 is where I was off at,

OpenStudy (amistre64):

000 to 099 is 100 excluded ways, + the 3 mentioned total 103 excluded ways 1000 103 ----- 897 ways total then sound right?

OpenStudy (amistre64):

im still including that 3rd digit ....

OpenStudy (amistre64):

00 to 99 is the last 2 digit options (100); exclude 00,01,02 (-3) 100-3 = 97 ways 6.9.97 = 5238

OpenStudy (anonymous):

thanks 5382 right? was that a typo error?